OpenStudy (cicilybailey14):

Fan+Medal Each pair of points is on the graph of an inverse variation. Find the missing value. (50, 13) and (t, 5) t = 25 t = 75 t = 110 t = 130

OpenStudy (cicilybailey14):

@Secret-Ninja @ganeshie8 ?

OpenStudy (solomonzelman):

well, the inverse variation looks like: \(\Large\color{black}{ y=\frac{k}{x} }\) so for example, if I had 2 points: (8,1) and (4,w) then to find w this is what I do: 1) Find the k, by plugging in the given point: \(\Large\color{black}{ 8=\frac{k}{1} }\) \(\Large\color{black}{ k=8 }\) 2) re-write the variation, and plug in the second point with a w. \(\Large\color{black}{ y=\frac{8}{x} }\) (re-wrote it now plugging in) \(\Large\color{black}{ 4=\frac{8}{w} }\) \(\Large\color{black}{ 4w=8 }\) \(\Large\color{black}{ w=2 }\)

OpenStudy (cicilybailey14):

I'm still thoroughly confused...But i'll manage lol. :D

OpenStudy (solomonzelman):

well: - write the equation for the inverse variation: - plug in your point (10,13) - solve for k - re-write the inverse variation but with the new value of k that you found. - then plug your point (t,5) into this (new) inverse variation, (with a found k) - solve for t